40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 26 17:07:41 2019, gautam, Update, LSC, POX / POY calibration POX_POY_sensorNoise.pdfALSnoise_20191125.pdf
    Reply  Tue Nov 26 18:16:08 2019, rana, Update, LSC, POX / POY calibration 
       Reply  Wed Nov 27 12:16:52 2019, gautam, Update, LSC, ITMX and ITMY OSEMs with low and high circulating power ITMXshadowSensors.pdfITMYshadowSensors.pdf
          Reply  Wed Nov 27 13:14:02 2019, rana, Update, LSC, ITMX and ITMY OSEMs with low and high circulating power 
             Reply  Wed Nov 27 16:10:29 2019, gautam, Update, LSC, AOM reconnected 
Message ID: 15053     Entry time: Wed Nov 27 16:10:29 2019     In reply to: 15052
Author: gautam 
Type: Update 
Category: LSC 
Subject: AOM reconnected 

i reconnected the AOM driver to the AOM in the main beam path (it was hijacked for the AOM in the AUX laser path for Anjali's MZ experiment). I also temporarily hooked up the AOM to a CDS channel to facilitate some swept-sine measurements. This was later disconnected. The swept sine will need some hardware to convert the bipolar drive signal from the CDS system to the unipolar input that the AOM driver wants (DTT swept sine wont let me set an offset for the excitation, although awggui can do this).

Quote:

if the RP don't fit

u must acquit

sweep the laser amplitude

to divine the couplin w certitude

ELOG V3.1.3-